When both copying machines A and B are working, 100 copies can be made in one minute. In one minute, copiers A and C can make 140 copies. If all three copy machines are working, 180 copies can be made in one minute. How many copies per minute can each copier make?
Answer by stanbon(48568)

When both copying machines A and B are working, 100 copies can be made in one minute. In one minute, copiers A and C can make 140 copies. If all three copy machines are working, 180 copies can be made in one minute. How many copies per minute can each copier make?
Let A, B, C be the number of copies produced respectively by the machines in one minute.
----------------
Each one's production rate is 1/A, !/B, !/C
----------------
A + B = 100
A + C = 140
A + B + C = 180
--------------------
(I used the matrix method on a TI-83 calculator to get:
A = 60 copies per minute
B = 40 copies per minute
C = 80 copies per minute
==================
Cheers,
Stan H.